Give the correct final term for the following expression to ensure that the resulting trinomial is a perfect square trinomial.
x^2+8x
Please show me the steps to get the answer. Thanks

Answer by Alan3354(69443)

x^2+8x
----------
a = 1, 1x^2
b = 8, 8x
-------------
b/2a = 8/2 = 4
c = (b/2a)^2 = 16
--> x^2 + 8x + 16 = (x + 4)^2